OpenStudy (bruno102):

Graph this f(x)=√x+2-3. I actually have an answer but don't know if it is correct.

OpenStudy (bruno102):

 mine was similar to this but more accurate on a graph

OpenStudy (anonymous):

not exactly... do you understand the shifts and such?

OpenStudy (bruno102):

Yes

OpenStudy (bruno102):

What I don't understand is how to change the equation to see the shifts (a, h, k)

OpenStudy (anonymous):

set whatever is inside the sqrt = 0 and solve. that will get you h. the k is whatever is outside is the k

OpenStudy (bruno102):

so I should solve √x+2

OpenStudy (anonymous):

set x+2 = 0 and solve for x. that will be your h.

OpenStudy (bruno102):

how am I supposed to solve if it is 0 I don't understand.

OpenStudy (anonymous):

subtract 2 from both sides to solve for x

OpenStudy (bruno102):

so f(x)=0-5?

OpenStudy (bruno102):

I mean f(x)= √0-5

OpenStudy (anonymous):

\[f(x) = \sqrt{x+2} - 3 \text{, yes or no?}\]

OpenStudy (bruno102):

yes

OpenStudy (anonymous):

so inside the square root, what value of x makes the inside = 0?

OpenStudy (bruno102):

-2

OpenStudy (anonymous):

great! that is h. in terms of x, is -2 to the left or right of the origin?

OpenStudy (bruno102):

left?

OpenStudy (anonymous):

correctamundo! now, that negative 3 is a shift in the y direction. is that going to be up or down?

OpenStudy (bruno102):

down I believe

OpenStudy (anonymous):

that's right! so your graph will pellet 2 to the left and down 3.

OpenStudy (anonymous):

shift not pellet

OpenStudy (bruno102):

How about the stretch and and reflection? (b and a)?

OpenStudy (anonymous):

there isn't any of those in this problem. if f(x) is the parent, then g(x) = af(x-h) + k is a horizontal shift of h, a vertical shift of k and a stretch of a if |a|>1 or a compression if |a| <1. if a< 0, then it is a reflection.

OpenStudy (bruno102):

see that is not what my teacher taught us, she said we have to factor the x+2 out of the equation to give us all of them but thanks.

OpenStudy (anonymous):

i'm not sure what your teacher is going for but i'm really not seeing the connection here. good luck